# A Statistical Test for Comparing the Linkage and Admixture Model Based on Central Limit Theorems

### v1

@@ Abstract
This study investigates the consistency and central limit behavior of maximum likelihood estimators (MLEs) for individual ancestry in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model. These theoretical results are used to prove theoretical properties of a statistical test that allows for model selection between the Admixture Model and the Linkage Model. Finally, we demonstrate the practical relevance of our results by applying the test to real-world data from The 1000 Genomes Project Consortium (2015).
@@ 1 Introduction
Both the Linkage Model (Falush et al., 2003) and the Admixture Model (e.g. Alexander et al. (2009)) are widely used to explain the genetic data of individuals. They are similar in that both assume that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations. However, the Linkage Model additionally incorporates the distance between loci through a Hidden Markov Model (HMM). Furthermore, in the Linkage Model, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event (Falush et al., 2003). As a result, Falush et al. (2003) formulated the Linkage Model as a Hidden Markov Model (HMM). The Admixture Model can be seen as a special case of the Linkage Model with $r=\infty$, i.e. the data is assumed to be independent across markers.
A natural question that arises is which model fits a given dataset better. This leads to a nested model selection problem as described by Anderson and Burnham (2004). In our case, we consider the statistical hypothesis test:
@@
This is a classical nested test problem, for which asymptotic theory exists (Wilks, 1938). However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry and $r$ are asymptotically normally distributed under the Linkage Model. Establishing asymptotic normality requires results from the theory of HMMs.
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al. (2005). However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
Several statistical tests and heuristics have been proposed for model selection in genetics. For example, to decide whether a pair of loci is suitable for STRUCTURE (Pritchard et al., 2000), Kaeuffer et al. (2007) suggested using $r_{LD}$, a measure of linkage disequilibrium introduced by Hill and Robertson (1968). They ran STRUCTURE with the Linkage Model and used logistic regression to assess the impact of $r_{LD}$ on the detection of population structure. Other studies have examined the sensitivity of cluster identification to genetic and ecological constraints (Rosenberg et al., 2001; Evanno et al., 2005).
To achieve our goal of establishing the theoretical properties of the test in (1), we proceed as follows: First, we precisely define both the Admixture Model and the Linkage Model. Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of chromosomes and when the number of markers tends to infinity. Based on these results, we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test. We also provide a method to quantify the uncertainty of the MLEs in the Linkage Model. Finally, we evaluate the statistical test through simulations and apply it to real data from The 1000 Genomes Project Consortium (2015).
@@ 2 Models
We first define the Linkage Model for one individual and bi-allelic markers.
@@
The number of alleles at chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ at chromosome $c$. Based on this, we define the Markov process
@@
We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{M_{C}},$
@@
We write $(q^{0},r^{0})$ for the true parameters $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
We always assume that $d_{c,m}\geq 0$ for all $m\in\{1,\ldots,M\}$ and that the allele frequencies and $d_{c,m}$ are known.
@@
Let $Z^{j}:=(Z^{j}_{c,1},...,Z^{j}_{c,M}),j=1,2,$ be two independent Markov chains with transition matrix (2). Let $X_{c,m}^{dip}$ be the number of alleles at chromosome $c,$ marker $m$ and let $x\in\{0,1,2\}.$ The emission probabilities for the diploid case are defined by
@@
While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems.
@@
The Admixture Model is a special case of the Linkage Model with $r=\infty.$
@@
We do not only need to know about the Linkage Model, but we also need some general notation is mentioned in Notation 2.4.
Notation 2.4 (General Notation).
For a positive definite matrix $A$, we write $A\succ 0.$ Additionally, we write $X_{i:i+s}$ for $X_{i},...,X_{i+s}.$
@@ 3 Theoretical Main Results
In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we need assumption 3.1.
@@
We assume that
  • (A1)
    the parameter space $\Theta:=\{[\kappa_{q},\kappa_{q}^{\prime}]^{K-1}\times[r_{lb},\infty]\}\subset\mathbb{R}^{K+1}$ is a compact set.
@@
    $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
@@
    Let $\lambda_{c,m}$ be the second largest of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0$
@@
ConsistencyCLTAsymptotic Distribution of the Test Statistic
Figure 1: Overview of results in this section and how they are connected to each other.
@@
We define the Fischer information
$J_{q^{0},r^{0}}:=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big{|}_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
and assume $J_{q^{0},r^{0}}\succ 0$. Let the true parameter be in the interior of the parameter space. Then, it holds
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J_{q^{0},r^{0}}^{-1}\right).$
@@
Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are all at different chromosomes or at the same chromosome.
According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 holds.
Definition 3.4 (Statistical Test).
We define the test statistic for the test (1) by
$\Lambda:=-2\ln\left(\frac{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,\infty))\}}{\textup{argmax}\{(q,r)\mapsto\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
@@ 4 Application to Data
In this chapter, we first evaluate the performance of the test from Definition 3.4 by using simulated data. Afterwards, we apply the test to data from The 1000 Genomes Project Consortium (2015) and compare the uncertainty of the MLEs in the Linkage Model to the ones in the Admixture Model.
@@ 4.1 Simulated Data
To evaluate the statistical test by using the type 1 and type 2 error, we simulate the data according to the Linkage Model and the Admixture Model with different values for $M,r$ and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}-1}.$ More precisely, we simulated 100 markers on one single chromosome with $d_{1,m}=d_{1,m+1}\forall m,d:=d_{1,m}\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}.$ We choose $0.05$ as the significance level. We calculated the type 1 and the type 2 error by repeating the experiment 100 times per possible combination of $r$ and $d$.
The results are shown in Figure 2. Especially the type 1 error is small and lower than 0.05. Furthermore, for increasing $r$ and $d$ the type 1 and the type 2 error tend to increase. This is not surprising since a large $r$ corresponds to a high similarity between the two models. To summarize, the simulation results show that the test has both, a high power and a small type 1 error.
Figure 2: Evaluation of the statistical test for simulated data.
@@ 4.2 Real Data
We also applied our theoretical results to the data from The 1000 Genomes Project Consortium (2015) with the AIM set by Kidd et al. (2014) that consists of $M_{total}=55$ markers on 20 different autosomes ($C=20$). We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency. The values for $d$ for the AIM set by Kidd et al. (2014) in cM are shown in Figure 3.
Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014). We rounded them to integer values.
The distance between the loci at a single chromosome was between below 1 cM and approximately 129 cM. We used individuals from Africa (AFR), Europe (EUR), South-East-Asian (SAS), East-Asian (EAS) and Admixed Americans (AMR), i.e. $K=5$.
The results of the statistical test are shown in Figure 4. Altogether, for 312 out of 2504 ($\approx 12.5\%$ of the) individuals, the null hypothesis that the Admixture Model fits better to the data, cannot be rejected. Thereby, we see strong differences between the populations: For people from Africa and East Asian, the fraction of individuals for which the null hypothesis is rejected is much smaller than for the other populations. This emphasizes that it is not sufficient to name a general rule to decide for either the Linkage Model or the Admixture Model that only depends on the marker set. Instead, a method – like the statistical test (1) – is required, that decides for every individual or population separately.
Figure 4: Results of the statistical test.
We also compare the covariance for both, the MLE in the Linkage Model and the MLE in the Admixture Model to each other for individual HG00096 from Great Britain. Therefore, we again used the AIM Set by Kidd et al. (2014). The results are presented in figures 6 and 6. There, we see that the variance of the estimators in the Linkage Model are higher than in the Admixture Model. This aligns with our expectations since the Linkage Model contains more parameters. Although both models clearly identify EUR as the primary ancestral population, the estimated values differ significantly.
@@
We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub. Therefore, we could either assume that $r$ depends on the individual, i.e. we would have $NK$ parameters that we have to estimate or that we have only one global parameter $r,$ i.e. we estimate N(K1)+1N(K-1)+1 parameters. We decided for the latter case. However, the allele frequencies are supposed to be known in this case. Additionally, the diploid case is also implemented.

### v2

# Highlights

@@ Abstract
We prove consistency and asymptotic normality of maximum likelihood estimators (MLEs) for the ancestry of individuals in the Linkage Model, complementing earlier results by (Pfaff et al., 2004; Pfaffelhuber and Rohde, 2022; Heinzel, 2025) for the Admixture Model. These results are used to prove that a statistical test that allows for model selection between the Admixture Model and the Linkage Model is an asymptotic level-$\alpha$-test. Finally, we demonstrate the practical relevance of our results by applying the test to real-world data from The 1000 Genomes Project Consortium (2015).
@@ 1 Introduction
The Linkage Model (Falush et al., 2003) is widely used to explain the genetic data of individuals. It assumes that genetic data can be described by the ancestry proportions $q$ of an individual from $K$ ancestral populations and the allele frequencies in these populations. Furthermore, there exists a parameter $r$, which can be interpreted as the number of generations since an admixture event. The genetic distance between loci is also considered by using a Hidden Markov Model (HMM). The Admixture Model can be seen as a special case of the Linkage Model with $r=\infty$, i.e. the data is assumed to be independent across markers.
A natural question that arises is which model, the Admixture Model or the Linkage Model, fits a given dataset better. This leads to a nested model selection problem as described by e.g. Anderson and Burnham (2004). In our case, we consider the statistical hypothesis test
@@
This is a classical test problem, for which asymptotic theory exists (Wilks, 1938). However, the test is only valid as an asymptotic level-$\alpha$ test if the MLEs for both ancestry $q$ and $r$ are asymptotically normally distributed under the Linkage Model. Proving asymptotic normality requires results from the theory of HMMs.
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al. (2005). However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution for the hidden chain across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
The question whether the data should be used in STRUCTURE (Pritchard et al., 2000)(a software that estimates the ancestry and the allele frequencies for the Admixture Model or the Linkage Model) analysis has already been considered. For example, to decide whether a pair of loci is suitable for STRUCTURE, Kaeuffer et al. (2007) suggested using $r_{LD}$, a measure of linkage disequilibrium introduced by Hill and Robertson (1968). They ran STRUCTURE with the Linkage Model and used logistic regression to assess the impact of $r_{LD}$ on the detection of population structure.
To achieve our goal of establishing the theoretical properties of the test in (1), we proceed as follows: First, we precisely define both the Admixture Model and the Linkage Model. Next, we prove the asymptotic normality of the MLEs in the Linkage Model, both when the number of observations tends to infinity. Based on these results, we quantify the uncertainty of the MLEs and we construct a statistical test for (1) and prove that it is an asymptotic level-$\alpha$ test. Finally, we evaluate the statistical test through simulations and apply it to real data from The 1000 Genomes Project Consortium (2015).
@@ 2 Models
We first define the Linkage Model for one individual and bi-allelic markers in Definition 2.1. The number of alleles on chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Recall that one centi Morgan corresponds to a $0.01$ crossing over between two genetic markers per Meiosis. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$. We write $(q^{0},r^{0})$ for the true values of $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
@@
We define the Markov chain
@@
We define the log-likelihood, for $M_{total}:=\sum_{c=1}^{C}{M_{c}},$
@@
We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}=1\}\subset\mathbb{R}^{K+1}$.
We always assume that the allele frequencies and $(d_{c,m})_{c=1,...,C,m=1,...,M_{c}}$ are known. Definition 2.1 is for haploid individuals, which can easily be extended to the diploid case.
@@
Let $Z^{j}:=(Z^{j}_{c,1},...,Z^{j}_{c,M}),j=1,2,$ be two independent Markov chains with transition matrix (2). Let $X_{c,m}^{dip}$ be the number of alleles on chromosome $c,$ marker $m,$ for diploid individuals. The emission probabilities for the diploid case are defined by
@@
While phased data is important for the Linkage Model, the Admixture Model can deal with unphased data without any problems. It is defined in Definition 2.3.
@@
The Admixture Model is a special case of the Linkage Model with $r=\infty.$ Especially, this means that it holds
$X_{c,m}\sim\textup{Bin}(2,\langle q_{\cdot}^{0},p_{c,\cdot,m}\rangle).$
@@
Finally, we define the statistical test that corresponds to test problem 2.
Definition 2.4 (Statistical Test).
We define the test statistic for the test (1) by
$\Lambda:=-2\ln\left(\frac{\max\left\{\ell((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
@@ 3 Theoretical Main Results
In this chapter, we prove the main results, i.e. consistency and central limit results, if the number of markers, $M_{total},$ tends to infinity. Therefore, we assume that Assumption 3.1 holds.
@@
We assume that $C=1$ and that we have haploid individuals. We additionally assume that
@@
    it holds $p_{k,m}\neq p_{\ell,m}$ for infinitely many markers m and for all $k,\ell\in\{1,..,K\},k\neq\ell.$
@@
    Let $\lambda_{c,m}$ be the second largest eigen value of the transition matrix $(\mathbb{P}(Z_{c,m}=k|Z_{c,m-1}=k^{\prime}))_{k=1,...,K,k^{\prime}=1,...,K}$. It holds
    $\prod_{c=1}^{C}\prod_{m=1}^{M_{c}}\lambda_{c,m}\xrightarrow[]{M_{total}\to\infty}0.$
@@
  • (A5)
    It holds $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$
By the definition of $\Theta$, we already assumed that $q\in[\kappa_{q},\kappa^{\prime}_{q}]$ and $r\geq r_{lb}.$
@@
Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible Fischer Information Asymptotic Distributionof the Test StatisticAssumption 3.1
Figure 1: Overview of results in this section and their requirements.
@@
Let Assumption 3.1 holdand let the Fischer information
$J(q^{0},r^{0}):=\lim_{M_{total}\to\infty}\mathbb{E}\left(\frac{\partial^{2}}{\partial(q,r)^{2}}\ell\left((x_{1,1},...,x_{C,M}),(q,r))\big|_{(q,r)=\left(q^{0},r^{0}\right)}\right)\right)$
be positive definite. Let the true parameters be in the interior of the parameter space. Then, it holds
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
@@
Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are either all on different chromosomes or on the same chromosome.
According to Wilks (1938), since the MLE in the Linkage Model is asymptotically normally distributed, Theorem 3 about the asymptotic distribution of the test statistic holds.
@@ 4 Application to Data
In this section, we first evaluate the performance of the test from Definition 2.4 by using simulated data. Afterwards, we apply the test to data from The 1000 Genomes Project Consortium (2015) and compare the uncertainty of the MLEs in the Linkage Model to the ones in the Admixture Model.
@@ 4.1 Simulated Data
To evaluate the statistical test in terms of Type 1 and Type 2 errors, I simulated data under both the Linkage Model and the Admixture Model for different values of $M$, $r$, and the marker distances $(d_{c,m})_{c=1,\dots,C,\ m=1,\dots,M_{c}-1}$. Specifically, I simulated 100 markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}$. The significance level was set to $0.05$.
I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$. The results are shown in Figure 2. Notably, the Type 1 error remains small and consistently below the $0.05$ threshold. Moreover, both the Type 1 and Type 2 errors tend to increase with increasing values of $r$ and $d$. This behavior is expected, as larger values of $r$ and $d$ imply greater similarity between the two models, making them harder to distinguish.
In summary, the simulation results indicate that the test exhibits high power while maintaining a low Type 1 error rate.
Figure 2: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$.
@@ 4.2 Real Data
I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with the ancestry informative marker (AIM) set by Kidd et al. (2014) consisting of $M_{total}=55$ markers on 20 different autosomes ($C=20$). For the analysis, individuals from five populations were included: Africa (AFR), Europe (EUR), South Asia (SAS), East Asia (EAS), and Admixed Americans (AMR), i.e., $K=5$. We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency. The values for $d$ for the AIM set by Kidd et al. (2014) in cM are shown in Figure 3.
Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014) rounded to integer values.
The distance between loci on a single chromosome ranged from below 1 cM to approximately 129 cM.
The results of the statistical test are shown in Figure 4. In total, for 312 out of 2504 individuals (approximately 12.5%), the null hypothesis — that the Admixture Model provides a better fit to the data — cannot be rejected. Notably, there are substantial differences across populations: for individuals from Africa and East Asia, the proportion for which the null hypothesis is rejected is considerably lower compared to other groups. This highlights that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set. Instead, a method — such as the statistical test defined in equation (1) — is needed, which allows model selection on an individual or population-specific basis.
Figure 4: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015).
We further compare the covariance matrices of the MLEs obtained under the Linkage Model and the Admixture Model for individual HG00096 from Great Britain. For this analysis, I again used the AIM set introduced by Kidd et al. (2014). The results are shown in Figure 6 and Figure 6, corresponding to the Admixture and Linkage Models, respectively.
As expected, the variance of the estimators is higher in the Linkage Model due to the larger number of parameters. While both models clearly identify EUR as the primary ancestral population, the estimated admixture proportions differ significantly between models (estimated ancestry of almost 1 from Europe in the Admixture Model and estimated ancestry of approximately 0.8 from Europe in the Linkage Model).
@@
We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub. In this framework, two modeling approaches are possible: either (i) the recombination rate $r$ is assumed to vary across individuals, requiring the estimation of $N\times K$ parameters, or (ii) a single global recombination rate $r$ is assumed, reducing the number of parameters to $N\times(K-1)+1$. We chose the latter approach (ii) to reduce model complexity.